OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 4.5 qt / min = ____ gal / h a.67.5 c.56.3 b.33.8 d.65

OpenStudy (anonymous):

@hilbertboy96

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

1 gal = 4qt and 1 h = 60 minutes. So, \(4.5 qt/min \times \cfrac{60 min.}{1 h} \times \cfrac{1 gal }{ 4qt} \) As , 60 min. / 1h = 1 and 1 gal / 4 qt = 1

mathslover (mathslover):

i.e. I just multiplied and divided by 1 : \(4. 5 \times \cfrac{1}{1} \times \cfrac{1}{1}\) Now as , 1/1 = 60 minute / 1 hour So, \(4.5 \times \cfrac{60 mn.}{1 h} \times \cfrac{1}{1}\) Also, 1/1 = 1 gal / 4qt So, \(4.5 qt / min. \times \cfrac{60 min. }{ 1 h.} \times \cfrac{1 gal}{ 4 qt}\) \(4.5 \cancel{qt} / \cancel{min.} \times \cfrac{60 \cancel{min.} }{1 h} \times \cfrac{1 gal}{4 \cancel{qt}}\)

mathslover (mathslover):

Solve further

OpenStudy (anonymous):

270?

mathslover (mathslover):

You forgot the "4" in the denominator \(4.5 \times \cfrac{60}{\boxed{4}}\)

OpenStudy (anonymous):

Lol. Yeah. Thank you. (:

mathslover (mathslover):

You're welcome @hadleydarcym :)